# Rank Two Sheaves With Low Discriminant on the Fano Threefold of Index 2 and Degree 5

### v1

@@ 2. The Derived Category $\mathrm{D}^{b}(X)$ and Stability
Proof.
All the statements are proved in [24, Lemma 6.2], except for the case of a strictly $\nu_{\alpha_{0},\beta_{0}}$-semistable object $E$. However, such an object also belongs to the category $\mathcal{A}^{{\alpha_{0}},{\beta_{0}}}(X)$ and has $\lambda_{{\alpha_{0}},{\beta_{0}},s}(E)=+\infty$, hence it is $\lambda_{{\alpha_{0}},{\beta_{0}},s}$-semistable. ∎
Recall that $X=X_{5}$ is a codimension 3 linear section of $\mathop{\mathrm{Gr}}\nolimits(2,V)$, where $V\cong\mathbb{C}^{5}$. Denote by $\mathcal{U}$ and $\mathcal{Q}$ the restrictions to $X$ of tautological subbundle and tautological quotient bundle on the Grassmannian, so that we have a short exact sequence
@@
In $\mathrm{D}^{b}(X)$ we have a full strong exceptional collection $(\mathcal{O}_{X}(-1),\mathcal{Q}(-1),\mathcal{U},\mathcal{O}_{X})$, it is the collection $\varepsilon_{2}$ from [20] up to a twist with a line bundle. More precisely, Orlov proved that it is a strong exceptional collection. Kuznetsov notes in [12, (12)] that $(\mathcal{U},\mathcal{Q}^{\vee},\mathcal{O}_{X},\mathcal{O}_{X}(1))$ is a full exceptional collection in $\mathrm{D}^{b}(X)$, and the collection that we use can be obtained from it by mutations and twisting with line bundles ([8, Lemma 5.3]), hence it is also full. The Chern classes of $\mathcal{U}$ and $\mathcal{Q}$ are calculated in [13]. Using this we can obtain the following lemma.
Lemma 2.11.
@@
(5) $D:=\{(\beta,\alpha)\in\mathbb{H}\ |\ \beta<-\frac{1}{2},\alpha<\beta+1\}.$
@@
Proof.
(i) We can plot the sets of points $(\beta,\alpha)\in\mathbb{R}^{2}$ where two of complex numbers $Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{O}_{X}(-1)),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{Q}(-1)),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{U}),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{O}_{X})$ are linearly dependent over $\mathbb{R}$.
Figure 1. Curves in which some of $Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{O}_{X}(-1)),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{Q}(-1)),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{U}),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{O}_{X})$ are $\mathbb{R}$-linearly dependent. The horizontal line is the $\beta$-axis, the vertical one is the $\alpha$-axis.
@@
where $\mathcal{U}^{\prime}\in\mathcal{T}^{\prime}_{\alpha,\beta},\mathcal{U}^{\prime\prime}\in\mathcal{F}^{\prime}_{\alpha,\beta}$. Then $\mathcal{U}^{\prime}[1]\in\mathcal{A}^{\alpha,\beta}(X)[1],\mathcal{U}^{\prime\prime}[1]\in\mathcal{A}^{\alpha,\beta}(X)$, so $\mathcal{U}[1]\in\nolinebreak\langle\mathcal{A}^{\alpha,\beta}(X),\mathcal{A}^{\alpha,\beta}(X)[1]\rangle$. ∎
@@
(6) $E^{\prime}\overset{a}{\to}E\to E^{\prime\prime}\to E^{\prime}[1]$
@@
Proof.
By [4, Proposition 2.3.2 (b)] $(\mathcal{A}\cap\mathcal{C},\mathcal{A}\cap(\mathcal{C}[-1]))$ is a torsion pair on the category $\mathcal{A}$. In particular, $\mathcal{A}\subset\langle\mathcal{C}[-1],\mathcal{C}\rangle$. Applying [4, Proposition 2.3.2 (b)] again, we get that $(\mathcal{C}[-1]\cap\mathcal{A},\mathcal{C}[-1]\cap(\mathcal{A}[-1]))$ is torsion pair on the category $\mathcal{C}[-1]$. Hence $(\mathcal{C}\cap(\mathcal{A}[1]),\mathcal{C}\cap\mathcal{A})$ is a torsion pair on the category $\mathcal{C}$ and for any $E\in\mathcal{C}$ we have an exact triple (6). ∎
@@
(7) $\begin{split}&\mathcal{T}^{\prime\prime}_{\gamma}=\{E\in\mathcal{A}^{\alpha,\beta}(X)\ |\ \text{any quotient}\ E\twoheadrightarrow G\,\text{satisfies}\,\lambda_{\alpha,\beta,s}(G)>\gamma\},\\ &\mathcal{F}^{\prime\prime}_{\gamma}=\{E\in\mathcal{A}^{\alpha,\beta}(X)\ |\ \text{any subobject}\ 0\neq F\hookrightarrow E\,\text{satisfies}\,\lambda_{\alpha,\beta,s}(F)\leq\gamma\}.\end{split}$
@@
Proof.
We take $\gamma=-\cot\pi\phi_{1}$, where $\phi_{1}$ satisfies the property (i) of Lemma 2.12. Since both $\mathcal{C}$ and $\langle\mathcal{T}^{\prime\prime}_{\gamma},\mathcal{F}^{\prime\prime}_{\gamma}[1]\rangle$ are hearts of bounded t-structures on $\mathrm{D}^{b}(X)$, it suffices to prove that $\mathcal{C}\subset\langle\mathcal{T}^{\prime\prime}_{\gamma},\mathcal{F}^{\prime\prime}_{\gamma}[1]\rangle$.
Take $E\in\mathcal{C}$ to show that $E\in\langle\mathcal{T}^{\prime\prime}_{\gamma},\mathcal{F}^{\prime\prime}_{\gamma}[1]\rangle$. We denote $\mathcal{A}:=\mathcal{A}^{\alpha,\beta}(X)$. By Lemma 2.13 we reduce the problem to the case when $E\in\mathcal{C}\cap\mathcal{A}$ or $E\in\mathcal{C}\cap(\mathcal{A}[1])$. Assume that $E\in\mathcal{C}\cap(\mathcal{A}[1])$ (for $E\in\mathcal{C}\cap\mathcal{A}$ one can give a dual argument). If $E\notin\mathcal{F}^{\prime\prime}_{\gamma}[1]$, then there exists a subobject $f:F\hookrightarrow E[-1]$ in $\mathcal{A}$ with $\lambda_{\alpha,\beta,s}(F)>\gamma$. Note that, since $0\neq F\in\mathcal{A}$ and $(\mathcal{A},Z_{\alpha,\beta,s})$ is a Bridgeland stability condition, we have $Z_{\alpha,\beta,s}(F)\neq 0$. By Lemma 2.13 there is an exact triangle $F^{\prime}\overset{f^{\prime}}{\to}F\to F^{\prime\prime}\to F^{\prime}[1]$ with $F^{\prime}\in\mathcal{A}\cap\mathcal{C},F^{\prime\prime}\in\mathcal{A}\cap(\mathcal{C}[-1])$. The object $F^{\prime}$ should be nonzero by the assumption on the $\lambda_{\alpha,\beta,s}$-slope of $F$ and the property (i) of Lemma 2.12, hence the morphism $f^{\prime}:F^{\prime}\to F$ is also nonzero. The composition $F^{\prime}\overset{f^{\prime}}{\to}F\overset{f}{\to}E[-1]$ has to be nonzero, since it is the composition of a nonzero morphism in $\mathcal{A}$ with a monomorphism $f$. However, $F^{\prime}\in\mathcal{C},E[-1]\in\mathcal{C}[-1]$, and existence of such a morphism contradicts the definition of a t-structure. ∎
@@
Also, for further reference we include the following result (the idea is from [7, Section 4]).
Lemma 2.17.
@@
Proof.
The result [11, Proposition 1.3] of Hartshorne together with Auslander-Buchsbaum formula implies that the projective dimension of $E$ at any point of $X$ is smaller than or equal to 1. Therefore, we can construct a resolution
@@
By [11, Corollary 1.4] the sheaf $E$ is locally free on the complement of a closed set of dimension $\leq 0$, hence $\dim\mathrm{supp}(\mathcal{E}xt^{1}(E,\mathcal{O}_{X}))\leq 0$. Such a sheaf can be obtained from skyscraper sheaves by extensions, and Grothendieck-Riemann-Roch Theorem implies, that $\mathrm{ch}_{i}(\mathcal{E}xt^{1}(E,\mathcal{O}_{X}))=0$ for $i\leq 2$, and $\mathrm{ch}_{3}(\mathcal{E}xt^{1}(E,\mathcal{O}_{X}))=h^{0}(\mathcal{E}xt^{1}(E,\mathcal{O}_{X}))$. Using additivity of Chern character in exact sequences and the property $\mathrm{ch}_{i}(\mathcal{E}^{\vee})=(-1)^{i}\mathrm{ch}_{i}(\mathcal{E})$ for a locally free sheaf $\mathcal{E}$, we obtain the statement of the lemma. ∎
3. Objects with Maximal Third Chern Character
@@
Proof.
We follow the lines of the proof of [24, Proposition 4.1]. Tensoring $E$ with $\mathcal{O}_{X}(-n)$ we reduce to the case $n=0$. We prove the statement for objects which are Bridgeland semistable for $\alpha=\frac{1}{4},\beta=-\frac{2}{3}$ and $s>\frac{1}{6}$ from Corollary 2.15. By this Corollary either $E$ or $E[1]$ belongs to $\mathfrak{C}$. The Chern character assumption implies that $E\in\mathfrak{C}$ and $E\cong\mathcal{O}_{X}^{\oplus m}$. The rest of the proof of the first statement repeats the arguments in [24, Proposition 4.1] verbatim.
The second statement is can analogously be reduced to the case $n=0$. Then the fact that $E\in\mathrm{Coh}^{\beta}(X)$ implies that $\beta<0$, and then the generalized Bogomolov–Gieseker inequality (Proposition 2.2) implies that $\mathrm{ch}_{3}(E)\leq 0$. ∎
@@
(8) $\widetilde{E}\to\mathbf{R}\mathcal{H}om(E,\mathcal{O}_{X})[1]\to T[-1]\to\widetilde{E}[1].$
Corollary 3.3.
@@
Proof.
Applying Proposition 3.2 to $E$ we obtain a tilt-semistable object $\widetilde{E}$ with $\mathrm{ch}_{\leq 2}(\widetilde{E})=\mathrm{ch}_{\leq 2}(\mathcal{O}_{X}(-n)^{\oplus m})$, and the exact triangle (8) gives $\mathrm{ch}_{3}(E)=\mathrm{ch}_{3}(\widetilde{E})-\mathrm{ch}_{3}(T)$. Proposition 3.1 says that $\mathrm{ch}_{3}(\widetilde{E})\leq\mathrm{ch}_{3}(\mathcal{O}_{X}(-n)^{\oplus m})$, which gives the desired bound on $\mathrm{ch}_{3}(E)$. In case when this bound is achieved we have $T=0$ and $\mathbf{R}\mathcal{H}om(E,\mathcal{O}_{X})[1]\cong\mathcal{O}_{X}(-n)^{\oplus m}$. Applying the dualizing functor $\mathbf{R}\mathcal{H}om(-,\mathcal{O}_{X})[1]$ to both sides we get $E\cong\mathcal{O}_{X}(n)^{\oplus m}[1]$. ∎
@@
Theorem 3.4.
Let $E\in\mathrm{Coh}^{\beta}(X)$ be an object with $\mathrm{ch}(E)=2+cH+dH^{2}+eH^{3}$, which is $\nu_{\alpha,\beta}$-semistable for all $(\beta,\alpha)$ in the closure $\overline{U}$ of a non-empty open subset $U\subset\mathbb{H}$.
@@
(1.2) If $d=-\frac{1}{10}$, then $e\leq\frac{7}{30}$. In case of equality $E$ is a coherent sheaf, included into an exact triple
$0\to\mathcal{U}(-1)\to\mathcal{O}_{X}(-1)^{\oplus 4}\to E\to 0.$
@@
(2.3) If $d=-\frac{2}{5}$, then $e\leq 0$. In case of equality $E$ is a coherent sheaf, included into an exact triple
$0\to\mathcal{Q}(-1)^{\oplus 2}\to\mathcal{U}^{\oplus 4}\to E\to 0.$
@@
Lemma 3.5.
If an object $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable in $\overline{U}$ for a non-empty open subset $U\subset\mathbb{H}$ and $\mathrm{ch}(E)=3-2H+\frac{3}{5}H^{2}+eH^{3},$ then $e\leq-\frac{7}{30}$.
Proof.
@@
Note that $E$ should be tilt-stable in $U$, since $\mathrm{ch}_{\leq 2}(E)\in\Lambda$ is a primitive element. If $E$ is not tilt-stable for some $\beta<-\frac{2}{3},\alpha>0$, then there is a semicircular wall $W$, which should intersect the ray $\beta=\beta_{-}(E)=\frac{-2-\sqrt{2/5}}{3}$ (by Proposition 2.3 (2, 3)). Lemma 2.9 implies that this wall is induced by a tilt-stable subobject or quotient $F$ with $\mathrm{ch}_{\leq 2}(F)=s+xH+yH^{2},s>0,\mu(F)<\mu(E)$. Since $0<\mathrm{ch}_{1}^{\beta_{-}(E)}(F)<\mathrm{ch}_{1}^{\beta_{-}(E)}(E)$, we get
(9) $0<x+\frac{2}{3}s+\frac{\sqrt{2/5}}{3}s<\sqrt{\frac{2}{5}}.$
@@
Also we have $1<\beta_{-}(E)<\beta_{-}(F)\leq\mu(F)<\mu(E)<0$ (cf. Proposition 2.4 (5)) and $\beta_{+}(F)=2\mu(F)-\beta_{-}(F)<2\mu(E)-\beta_{-}(E)<0$. Hence $\beta_{-}(F),\beta_{+}(F)\in[-1,0)$. If $s\neq 1$, we can apply Theorem 2.8 to $F$ and get that $s^{2}\leq\overline{\Delta}_{H}(F)<\overline{\Delta}_{H}(E)=10$. Hence, $s\in\{1,2,3\}$. A direct calculation shows that for any such $s$ there are no integer $x$ satisfying (9). Therefore, $E$ is tilt-stable for all $\beta<-\frac{2}{3},\alpha>0$.
@@
Proposition 2.7 implies that $E$ should be a 2-stable sheaf, in particular, $H^{0}(E)=0$. If $H^{2}(E)\neq 0$, then by Serre duality there is a nonzero morphism $E\to\mathcal{O}_{X}(-2)[1]$. However, such a morphism would destabilize $E$ below the wall $W(E,\mathcal{O}_{X}(-2)[1])$, contradicting what we have proved. We have $\mathrm{td}(T_{X})=1+H+\frac{8}{15}H^{2}+\frac{1}{5}H^{3}$ (see [21, Lemma 2.1]). Grothendieck-Riemann-Roch Theorem gives that $\chi(E)=5e+\frac{2}{3}=-h^{1}(E)-h^{3}(E)\leq 0$, hence $e\leq-\frac{2}{15}$.
@@
Since $1<\beta_{-}(E)<-\frac{1}{2}$, Lemma 2.10 implies that $E[1]$ is $\lambda_{\alpha,\beta,s}$-semistable for some $(\beta,\alpha)\in D$ with infinite $\lambda_{\alpha,\beta,s}(E[1])$. Then Corollary 2.15 implies that $E[1]$ belongs to $\mathfrak{C}$. For the case $e=-\frac{2}{15}$ we find that $E\cong\mathcal{O}_{X}(-1)[2]\oplus\mathcal{U}$. However, this object does not belong to $\mathrm{Coh}^{\beta}(X)$, so this case is impossible. Since the step by which $\mathrm{ch}_{3}$ can change for a fixed $\mathrm{ch}_{\leq 2}$ is equal to $\frac{H^{3}}{10}$ (because $c_{3}$ belongs to $H^{6}(X,\mathbb{Z})$), we get that $e\leq-\frac{7}{30}$. ∎
Lemma 3.6.
@@
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable in $\overline{U}$ for a non-empty open subset $U\subset\mathbb{H}$ and $\mathrm{ch}(E)=2-H+\frac{H^{2}}{10}+eH^{3}$, then $e\leq\frac{1}{30}$, and in case of equality we have $E\cong\mathcal{U}$.
Proof.

### v2

@@ 2 The Derived Category $\mathrm{D}^{b}(X)$ and Stability
Proof. All the statements are proved in [24, Lemma 6.2], except for the case of a strictly $\nu_{\alpha_{0},\beta_{0}}$-semistable object $E$. However, such an object also belongs to the category $\mathcal{A}^{{\alpha_{0}},{\beta_{0}}}(X)$ and has $\lambda_{{\alpha_{0}},{\beta_{0}},s}(E)=+\infty$, hence it is $\lambda_{{\alpha_{0}},{\beta_{0}},s}$-semistable. $\square$
Recall that $X=X_{5}$ is a codimension 3 linear section of $\mathrm{Gr}(2,V)$, where $V\cong\mathbb{C}^{5}$. Denote by $\mathcal{U}$ and $\mathcal{Q}$ the restrictions to $X$ of tautological subbundle and tautological quotient bundle on the Grassmannian, so that we have a short exact sequence
@@
In $\mathrm{D}^{b}(X)$ we have a full strong exceptional collection $(\mathcal{O}_{X}(-1),\mathcal{Q}(-1),\mathcal{U},\mathcal{O}_{X})$, it is the collection $\varepsilon_{2}$ from [20] up to a twist with a line bundle. More precisely, Orlov proved that it is a strong exceptional collection. Kuznetsov notes in [12, (12)] that $(\mathcal{U},\mathcal{Q}^{\vee},\mathcal{O}_{X},\mathcal{O}_{X}(1))$ is a full exceptional collection in $\mathrm{D}^{b}(X)$, and the collection that we use can be obtained from it by mutations and twisting with line bundles ([7, Lemma 5.3]), hence it is also full. The Chern classes of $\mathcal{U}$ and $\mathcal{Q}$ are calculated in [13]. Using this we can obtain the following lemma.
Lemma 2.11
@@
$D:=\{(\beta,\alpha)\in\mathbb{H}\ |\ \beta<-\frac{1}{2},\alpha<\beta+1\}.$ (6)
@@
Proof. (i) We can plot the sets of points $(\beta,\alpha)\in\mathbb{R}^{2}$ where two of complex numbers $Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{O}_{X}(-1)),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{Q}(-1)),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{U}),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{O}_{X})$ are linearly dependent over $\mathbb{R}$.
Рис. 1: Curves in which some of $Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{O}_{X}(-1)),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{Q}(-1)),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{U}),Z_{\alpha,\beta,\frac{1}{6}}(\mathcal{O}_{X})$ are $\mathbb{R}$-linearly dependent. The horizontal line is the $\beta$-axis, the vertical one is the $\alpha$-axis.
@@
where $\mathcal{U}^{\prime}\in\mathcal{T}^{\prime}_{\alpha,\beta},\mathcal{U}^{\prime\prime}\in\mathcal{F}^{\prime}_{\alpha,\beta}$. Then $\mathcal{U}^{\prime}[1]\in\mathcal{A}^{\alpha,\beta}(X)[1],\mathcal{U}^{\prime\prime}[1]\in\mathcal{A}^{\alpha,\beta}(X)$, so $\mathcal{U}[1]\in\nolinebreak\langle\mathcal{A}^{\alpha,\beta}(X),\mathcal{A}^{\alpha,\beta}(X)[1]\rangle$. $\square$
@@
$E^{\prime}\overset{a}{\to}E\to E^{\prime\prime}\to E^{\prime}[1]$ (7)
@@
Proof. By [3, Proposition 2.3.2 (b)] $(\mathcal{A}\cap\mathcal{C},\mathcal{A}\cap(\mathcal{C}[-1]))$ is a torsion pair on the category $\mathcal{A}$. In particular, $\mathcal{A}\subset\langle\mathcal{C}[-1],\mathcal{C}\rangle$. Applying [3, Proposition 2.3.2 (b)] again, we get that $(\mathcal{C}[-1]\cap\mathcal{A},\mathcal{C}[-1]\cap(\mathcal{A}[-1]))$ is torsion pair on the category $\mathcal{C}[-1]$. Hence $(\mathcal{C}\cap(\mathcal{A}[1]),\mathcal{C}\cap\mathcal{A})$ is a torsion pair on the category $\mathcal{C}$ and for any $E\in\mathcal{C}$ we have an exact triple (7). $\square$
@@
$\begin{split}&\mathcal{T}^{\prime\prime}_{\gamma}=\{E\in\mathcal{A}^{\alpha,\beta}(X)\ |\ \text{any quotient}\ E\twoheadrightarrow G\,\text{satisfies}\,\lambda_{\alpha,\beta,s}(G)>\gamma\},\\ &\mathcal{F}^{\prime\prime}_{\gamma}=\{E\in\mathcal{A}^{\alpha,\beta}(X)\ |\ \text{any subobject}\ 0\neq F\hookrightarrow E\,\text{satisfies}\,\lambda_{\alpha,\beta,s}(F)\leq\gamma\}.\end{split}$ (8)
@@
Proof. We take $\gamma=-\cot\pi\phi_{1}$, where $\phi_{1}$ satisfies the property (i) of Lemma 2.12. Since both $\mathcal{C}$ and $\langle\mathcal{T}^{\prime\prime}_{\gamma},\mathcal{F}^{\prime\prime}_{\gamma}[1]\rangle$ are hearts of bounded t-structures on $\mathrm{D}^{b}(X)$, it suffices to prove that $\mathcal{C}\subset\langle\mathcal{T}^{\prime\prime}_{\gamma},\mathcal{F}^{\prime\prime}_{\gamma}[1]\rangle$.
Take $E\in\mathcal{C}$ to show that $E\in\langle\mathcal{T}^{\prime\prime}_{\gamma},\mathcal{F}^{\prime\prime}_{\gamma}[1]\rangle$. We denote $\mathcal{A}:=\mathcal{A}^{\alpha,\beta}(X)$. By Lemma 2.13 we reduce the problem to the case when $E\in\mathcal{C}\cap\mathcal{A}$ or $E\in\mathcal{C}\cap(\mathcal{A}[1])$. Assume that $E\in\mathcal{C}\cap(\mathcal{A}[1])$ (for $E\in\mathcal{C}\cap\mathcal{A}$ one can give a dual argument). If $E\notin\mathcal{F}^{\prime\prime}_{\gamma}[1]$, then there exists a subobject $f:F\hookrightarrow E[-1]$ in $\mathcal{A}$ with $\lambda_{\alpha,\beta,s}(F)>\gamma$. Note that, since $0\neq F\in\mathcal{A}$ and $(\mathcal{A},Z_{\alpha,\beta,s})$ is a Bridgeland stability condition, we have $Z_{\alpha,\beta,s}(F)\neq 0$. By Lemma 2.13 there is an exact triangle $F^{\prime}\overset{f^{\prime}}{\to}F\to F^{\prime\prime}\to F^{\prime}[1]$ with $F^{\prime}\in\mathcal{A}\cap\mathcal{C},F^{\prime\prime}\in\mathcal{A}\cap(\mathcal{C}[-1])$. The object $F^{\prime}$ should be nonzero by the assumption on the $\lambda_{\alpha,\beta,s}$-slope of $F$ and the property (i) of Lemma 2.12, hence the morphism $f^{\prime}:F^{\prime}\to F$ is also nonzero. The composition $F^{\prime}\overset{f^{\prime}}{\to}F\overset{f}{\to}E[-1]$ has to be nonzero, since it is the composition of a nonzero morphism in $\mathcal{A}$ with a monomorphism $f$. However, $F^{\prime}\in\mathcal{C},E[-1]\in\mathcal{C}[-1]$, and existence of such a morphism contradicts the definition of a t-structure. $\square$
@@
Also, for further reference we include the following result (the idea is from [6, Section 4]).
Lemma 2.17
@@
Proof. The result [10, Proposition 1.3] of Hartshorne together with Auslander-Buchsbaum formula implies that the projective dimension of $E$ at any point of $X$ is smaller than or equal to 1. Therefore, we can construct a resolution
@@
By [10, Corollary 1.4] the sheaf $E$ is locally free on the complement of a closed set of dimension $\leq 0$, hence $\dim\mathrm{supp}(\mathcal{E}xt^{1}(E,\mathcal{O}_{X}))\leq 0$. Such a sheaf can be obtained from skyscraper sheaves by extensions, and Grothendieck-Riemann-Roch Theorem implies, that $\mathrm{ch}_{i}(\mathcal{E}xt^{1}(E,\mathcal{O}_{X}))=0$ for $i\leq 2$, and $\mathrm{ch}_{3}(\mathcal{E}xt^{1}(E,\mathcal{O}_{X}))=h^{0}(\mathcal{E}xt^{1}(E,\mathcal{O}_{X}))$. Using additivity of Chern character in exact sequences and the property $\mathrm{ch}_{i}(\mathcal{E}^{\vee})=(-1)^{i}\mathrm{ch}_{i}(\mathcal{E})$ for a locally free sheaf $\mathcal{E}$, we obtain the statement of the lemma. $\square$
3 Objects with Maximal Third Chern Character
@@
Proof. We follow the lines of the proof of [24, Proposition 4.1]. Tensoring $E$ with $\mathcal{O}_{X}(-n)$ we reduce to the case $n=0$. We prove the statement for objects which are Bridgeland semistable for $\alpha=\frac{1}{4},\beta=-\frac{2}{3}$ and $s>\frac{1}{6}$ from Corollary 2.15. By this Corollary either $E$ or $E[1]$ belongs to $\mathfrak{C}$. The Chern character assumption implies that $E\in\mathfrak{C}$ and $E\cong\mathcal{O}_{X}^{\oplus m}$. The rest of the proof of the first statement repeats the arguments in [24, Proposition 4.1] verbatim.
The second statement is can analogously be reduced to the case $n=0$. Then the fact that $E\in\mathrm{Coh}^{\beta}(X)$ implies that $\beta<0$, and then the generalized Bogomolov–Gieseker inequality (Proposition 2.2) implies that $\mathrm{ch}_{3}(E)\leq 0$. $\square$
@@
$\widetilde{E}\to\mathbf{R}\mathcal{H}om(E,\mathcal{O}_{X})[1]\to T[-1]\to\widetilde{E}[1].$ (9)
Corollary 3.3
@@
Proof. Applying Proposition 3.2 to $E$ we obtain a tilt-semistable object $\widetilde{E}$ with $\mathrm{ch}_{\leq 2}(\widetilde{E})=\mathrm{ch}_{\leq 2}(\mathcal{O}_{X}(-n)^{\oplus m})$, and the exact triangle (9) gives $\mathrm{ch}_{3}(E)=\mathrm{ch}_{3}(\widetilde{E})-\mathrm{ch}_{3}(T)$. Proposition 3.1 says that $\mathrm{ch}_{3}(\widetilde{E})\leq\mathrm{ch}_{3}(\mathcal{O}_{X}(-n)^{\oplus m})$, which gives the desired bound on $\mathrm{ch}_{3}(E)$. In case when this bound is achieved we have $T=0$ and $\mathbf{R}\mathcal{H}om(E,\mathcal{O}_{X})[1]\cong\mathcal{O}_{X}(-n)^{\oplus m}$. Applying the dualizing functor $\mathbf{R}\mathcal{H}om(-,\mathcal{O}_{X})[1]$ to both sides we get $E\cong\mathcal{O}_{X}(n)^{\oplus m}[1]$. $\square$
@@
Theorem 3.4
Let $E\in\mathrm{Coh}^{\beta}(X)$ be a tilt-semistable object with $\mathrm{ch}(E)=2+cH+dH^{2}+eH^{3}$.
@@
(1.2) If $d=-\frac{1}{10}$, then $e\leq\frac{7}{30}$. In case of equality $E$ is included into an exact triple
$0\to\mathcal{O}_{X}(-1)^{\oplus 4}\to E\to\mathcal{U}(-1)[1]\to 0.$
@@
(2.3) If $d=-\frac{2}{5}$, then $e\leq 0$. In case of equality $E$ is included into an exact triple
$0\to\mathcal{U}^{\oplus 4}\to E\to\mathcal{Q}(-1)^{\oplus 2}[1]\to 0.$
@@
Lemma 3.5
If an object $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable and $\mathrm{ch}(E)=3-2H+\frac{3}{5}H^{2}+eH^{3},$ then $e\leq-\frac{7}{30}$.
@@
Proof. If $E$ is not tilt-semistable for some $\beta<-\frac{2}{3},\alpha>0$, then there is a semicircular wall $W$, which should intersect the ray $\beta=\beta_{-}(E)=\frac{-2-\sqrt{2/5}}{3}$ (by Proposition 2.3 (2, 3)). Lemma 2.9 implies that this wall is induced by a tilt-stable subobject or quotient $F$ with $\mathrm{ch}_{\leq 2}(F)=s+xH+yH^{2},s>0,\mu(F)<\mu(E)$. Since $0<\mathrm{ch}_{1}^{\beta_{-}(E)}(F)<\mathrm{ch}_{1}^{\beta_{-}(E)}(E)$, we get
$0<x+\frac{2}{3}s+\frac{\sqrt{2/5}}{3}s<\sqrt{\frac{2}{5}}.$ (10)
@@
Also we have $1<\beta_{-}(E)<\beta_{-}(F)\leq\mu(F)<\mu(E)<0$ (cf. Proposition 2.4 (5)) and $\beta_{+}(F)=2\mu(F)-\beta_{-}(F)<2\mu(E)-\beta_{-}(E)<0$. Hence $\beta_{-}(F),\beta_{+}(F)\in[-1,0)$. If $s\neq 1$, we can apply Theorem 2.8 to $F$ and get that $s^{2}\leq\overline{\Delta}_{H}(F)<\overline{\Delta}_{H}(E)=10$. Hence, $s\in\{1,2,3\}$. A direct calculation shows that for any such $s$ there are no integer $x$ satisfying (10). Therefore, $E$ is tilt-semistable for all $\beta<-\frac{2}{3},\alpha>0$.
@@
Proposition 2.7 implies that $E$ should be a 2-semistable sheaf, in particular, $H^{0}(E)=0$. If $H^{2}(E)\neq 0$, then by Serre duality there is a nonzero morphism $E\to\mathcal{O}_{X}(-2)[1]$. However, such a morphism would destabilize $E$ below the wall $W(E,\mathcal{O}_{X}(-2)[1])$, contradicting what we have proved. We have $\mathrm{td}(T_{X})=1+H+\frac{8}{15}H^{2}+\frac{1}{5}H^{3}$ (see [21, Lemma 2.1]). Grothendieck-Riemann-Roch Theorem gives that $\chi(E)=5e+\frac{2}{3}=-h^{1}(E)-h^{3}(E)\leq 0$, hence $e\leq-\frac{2}{15}$.
@@
Since $1<\beta_{-}(E)<-\frac{1}{2}$, Lemma 2.10 implies that $E[1]$ is $\lambda_{\alpha,\beta,s}$-semistable for some $(\beta,\alpha)\in D$ with infinite $\lambda_{\alpha,\beta,s}(E[1])$. Then Corollary 2.15 implies that $E[1]$ belongs to $\mathfrak{C}$. For the case $e=-\frac{2}{15}$ we find that $E\cong\mathcal{O}_{X}(-1)[2]\oplus\mathcal{U}$. However, this object does not belong to $\mathrm{Coh}^{\beta}(X)$, so this case is impossible. Since the step by which $\mathrm{ch}_{3}$ can change for a fixed $\mathrm{ch}_{\leq 2}$ is equal to $\frac{H^{3}}{10}$ (because $c_{3}$ belongs to $H^{6}(X,\mathbb{Z})$), we get that $e\leq-\frac{7}{30}$. $\square$
Lemma 3.6
@@
If $E\in\mathrm{Coh}^{\beta}(X)$ is tilt-semistable and $\mathrm{ch}(E)=2-H+\frac{H^{2}}{10}+eH^{3}$, then $e\leq\frac{1}{30}$, and in case of equality we have $E\cong\mathcal{U}$.